A manufacturer knows that their items have a lengths that are skewed right, with a mean of 14 inches, and standard deviation of 3.1 inches. If 35 items are chosen at random, what is the probability that their mean length is greater than 14.4 inches? (Round answer to four decimal places)
Answer #1

Solution :

Given that ,

mean = = 14

standard deviation = = 3.1

n = 35

= = 14 and

= / n = 3.1 / 35 = 0.5240

P( > 14.4) = 1 - P( < 14.4)

= 1 - P(( - ) / < (14.4 - 14) / 0.5240)

= 1 - P(z < 0.76)

= 1 - 0.7764 Using standard normal table.

= 0.2236

The probability that their mean length is greater than 14.4 inches is 0.2236